# Shadows of the Colossus: Hierarchical Black Hole Mergers in a 10-million-body Globular Cluster Simulation

### v1

@@ 3.2 Black hole mergers
The $243.7~M_{\odot}$ fifth-generation black hole is particularly noted for its similarity in mass to the recent event GW231123, which has total mass $225^{+26}_{-43}~M_{\odot}$ (The LIGO Scientific Collaboration et al., 2025). Recent work in Passenger et al. (2025) placed low probability on a 2G+2G merger being responsible for GW231123, and we find this conclusion consistent with both colossus and lower-mass models. The highest-mass 2G+2G remnant in colossus, being $153.5~M_{\odot}$, lies well below the total mass of GW231123. On the other hand, the 3G+ mergers exclusive to GCs as massive as colossus are more comparable to the mass of GW231123. This suggests that the most massive GCs (comparable to masses of nuclear star clusters) indeed provide a channel for high-mass GW observations, via 3G+ mergers (e.g. Antonini & Rasio, 2016; Fragione & Rasio, 2023; The LIGO Scientific Collaboration et al., 2025). However, from the perspective of spins, we find the highest-mass mergers formed via 3G+ mergers in colossus generally have spin values lower than the $\chi_{1}=0.90^{+0.10}_{-0.19}$ and $\chi_{2}=0.80^{+0.20}_{-0.51}$ values quoted for GW231123 at $90\%$ confidence (for reasons described in Figure 4). In this case, spins may pose a challenge for formation of GW231123-like events via 3G+ mergers (for further discussion, see Passenger et al., 2025). Alternative formation channels that may produce high spins include accretion in active galactic nuclei disks (e.g., Tagawa et al., 2020; Vajpeyi et al., 2022; Bartos & Haiman, 2025) or from stars in dense clusters (e.g., Kıroğlu et al., 2025b), chemically homogeneous evolution (e.g, Marchant et al., 2016, 2024; Stegmann et al., 2025), and tidal spin up in stellar binaries (e.g., Bavera et al., 2021; Ma & Fuller, 2023; Qin et al., 2023).
@@
In the left panel of Figure 6 we show the delay time distribution for all mergers in colossus compared to low-mass CMC models (see Figure 2 and Section 4). As shown, more massive clusters generally feature longer delay times, simply because more massive clusters have longer relaxation times ($t_{\rm rlx}\propto M_{\rm cl}^{1/2}$; e.g., Binney & Tremaine, 2008). However at late times ($t\gtrsim 8\,$Gyr), we see the delay time distribution for colossus steepens relative to lower-mass models. This occurs because unlike lower-mass models, nearly all of the mergers in colossus occur inside the cluster, which in general have shorter GW inspiral times relative to ejected mergers (e.g., Rodriguez et al., 2018; Kremer, 2026). We address this point further in Section 4.
In Table 1, we list the number of BBH mergers of various types in the colossus model. “Ejected” mergers refer to BBHs that merge following dynamical ejection from the cluster, “2-body insp.” mergers refer to BBHs that merge inside their host cluster following dynamical hardening and GW inspiral, and “2-body cap.”, “3-body cap.”, and “4-body cap.” refer to in-cluster mergers that occur via GW capture during single-single, binary-single, and binary-binary encounters respectively. For further discussion of the distinction between these various channels, see Kremer et al. (2020b).
@@ 3.3 Colossus versus Behemoth
The most comparable CMC model run to-date is the behemoth model of Rodriguez et al. (2020), which was computed as part of the “Great Balls of FIRE” project designed to perform cluster simulations with initial conditions and time-dependent tidal forces extracted from a cosmological simulation (Grudić et al., 2023; Rodriguez et al., 2023; Bruel et al., 2024). The behemoth model also contained roughly $10^{7}$ stars at birth, but featured two key differences relative to our colossus model: (1) behemoth was run for only 44\,Gyr before it was fully disrupted by external tidal forces in its host galaxy and (2) behemoth is higher metallicity (roughly $0.5\,Z_{\odot}$), which translates to much lower mass black holes due to the influence of metallicity-dependent stellar mass loss due to winds (Vink et al., 2001). For example, the median primary mass for all first-generation black hole mergers in behemoth is $16.7\,M_{\odot}$, while it is $30.3\,M_{\odot}$ for colossus. From a numerical perspective, clusters with higher-mass black holes are computationally slower as the black holes undergo deeper collapse episodes driven by enhanced mass segregation (e.g., Kremer et al., 2020b; Rodriguez et al., 2022). This in turn leads to shorter time steps compared to higher-metallicity simulations, which results in longer run times. Nonetheless, the behemoth model features similar conclusions to those drawn from our new colossus model, namely that the vast majority of BBH mergers occur inside the cluster and that extended chains of hierarchical mergers are enabled by the relatively higher cluster escape velocity (for further detail and discussion of comparison to specific LVK events, see Rodriguez et al., 2020).
@@ 3.4 Other transient events
Although we focus here primarily on BBH mergers, for completeness, we also include in Table 1 the number of other types of mergers and collisions occuring in the colossus simulation. Although roughly 6,0006{,}000 neutron stars are formed and retained in the cluster via stellar evolution, zero binary neutron star mergers occur throughout the simulation. This is consistent with findings of previous studies (notably Ye et al., 2019) that have shown binary neutron star mergers are quite rare in clusters that have not yet reached cluster core collapse. We identify $404$ total white dwarf binary mergers in the simulation, all of which merge inside the cluster. This is also consistent with expectations from previous studies (e.g., Kremer et al., 2021) that have shown (1) white dwarf mergers preferentially occur inside their host cluster relative to BBH mergers, owing to their relatively small masses and correspondingly small semi-major axis values required for dynamical ejection and (2) similar to neutron star mergers, white dwarf mergers are rare relative to BH mergers in non-core-collapsed clusters.
We also list at the bottom of Table 1 counts of other types of collisions. Notably, we find $759$ BH+star collisions occur over in the complete simulation. For comparison, $85$ and $37$ BH+star collisions occur in the corresponding CMC Cluster Catalog models with $N=32\times 10^{5},r_{v}=2\,$pc and $N=16\times 10^{5},r_{v}=2\,$pc, respectively. These BH+star collisions may give rise to bright electromagnetic transients powered by accretion of the disrupted star onto the BH (Kremer et al., 2019b, 2022, 2023; Kıroğlu et al., 2023), potentially similar to transients events observed recently in the outskirts of nearby old massive elliptical galaxies (Nicholl et al., 2023). Similar to BBH mergers, we find there is a strong preference for BH+star collisions to occur in the most massive GCs, which may aid in potential follow-up of observed events as more massive GCs are brighter.
@@ 4.1 Scaling of BBH mergers with cluster mass
For comparison to colossus, we utilize the same five lower-mass CMC Cluster Catalog models plotted in previous figures, with the same initial conditions except for initial cluster $N$. In Figure 7, we plot the number of BBH mergers in all six models versus initial cluster mass, $M_{\rm cl,0}$, and initial $N$, sorted by merger channel, hierarchical generation, and location. In the left panel of Figure 7, we plot the total number of BBH mergers as a solid black curve, alongside three separate curves for the distinct merger channels (for review, see Kremer, 2026): ejected mergers (blue diamonds), in-cluster 2-body inspirals (orange squares), and in-cluster GW capture mergers (red triangles). Note that GW captures here combines single-single, binary-single, and binary-binary capture channels listed in Table 1. A linear regression fit to the total BBH curve in log-log space yields the best-fit relation
@@
For comparison, Antonini & Gieles (2020) uses semi-analytic methods to find $N_{\rm bbh}\propto M_{\rm cl,0}^{1.6}$ for BBH mergers occurring at late times (t>8t>8\,Gyr). Hong et al. (2018) used Monte Carlo models performed with the MOCCA code (Giersz et al., 2013) to find $N_{\rm bbh}\propto M_{\rm cl,0}^{1.3}$ for mergers occurring across all times. The $N_{\rm bbh}\propto M_{\rm cl,0}^{1.42}$ scaling found here is roughly consistent with these previous studies. We also performed a similar regression for present-day cluster mass $M_{\rm cl,f}$ (see Figure 2) and found a scaling of $N_{\rm bbh}\propto M_{\rm cl,f}^{1.36}$. Since the scaling with initial and final cluster masses are very similar, henceforth we simply adopt $N_{\rm bbh}\propto M_{\rm cl}^{1.4}$. Regression fits to other CMC models with varying $r_{v}$ show this scaling relation remains consistent, independent of $r_{v}$.
@@ 4.2 Scaling of BBH mergers with half-light radius
Similar to cluster mass, BBH merger populations are also expected to vary with cluster radius (e.g. Kremer et al., 2020b). We utilize 80 CMC models of varying initial $N$ and $r_{v}$ to derive a scaling relation in present-day cluster $r_{h}$, but exclude colossus in this analysis since no other model has comparable mass.
@@
As motivated by previous studies (Hong et al., 2018; Antonini & Gieles, 2020), we initially attempted to fit a power-law correlation to $r_{h}$ of the form $N_{\rm bbh}\propto r_{h}^{-\alpha}$, but found that power-law fits were unable to produce accurate results for the $r_{v}=0.5~\text{pc}$ CMC models (previous studies did not consider $r_{v}<1~\text{pc}$ clusters in their analyses). We were able to reproduce Antonini & Gieles (2020)’s $N_{\rm bbh}\propto r_{h}^{-0.9}$ relation by excluding the $r_{v}=0.5~\text{pc}$ models from the regression fit, yielding $N_{\rm bbh}\propto r_{h}^{-0.97}$ (plotted as a dashed curved in Figure 8), but this fit significantly overestimates $N_{\rm bbh}$ at low $r_{h}$. We instead utilize an exponential function of the form $N_{\rm bbh}\propto e^{-\alpha r_{h}}$ to better fit the data. The best-fit curves for the total BBH merger count with varying $r_{h}$ obtained through exponential regression are given in Equation 5, and plotted in Figure 8. The $N=2$ and $N=4$ models did not exhibit sufficiently distinct BBH merger populations, so they were grouped into one curve.
@@ 4.3 Ejected BBH properties
In the dense environment of GCs, BBHs experience repeated close encounters with other single stars and binary systems. Each of these encounters transfer energy from the BBH’s internal energy into the kinetic energy of the interacting objects, which leads to BBH hardening and dynamical recoil (e.g., Rodriguez et al., 2016; Kremer, 2026). Energetics arguments show center-of-mass recoil velocity of a binary is comparable to its orbital velocity $v_{\rm orb}\approx\sqrt{Gm_{\rm bh}/a}$. Therefore, as a BBH hardens, it recoils more. The characteristic minimum separation a binary can achieve via dynamical hardening is set by its host cluster’s escape velocity (Equation 1). Once a BBH becomes sufficiently compact, its recoil velocity will exceed its host’s escape velocity, and it will be ejected from the cluster. A subset of these ejected binaries will have sufficiently short GW inspiral times to ultimately merge within a Hubble time. Since higher-mass GCs have higher escape energy thresholds, ejected BBH systems of a higher-mass GC will have higher binding energies, shorter semi-major axes, and correspondingly shorter GW inspiral times (Rodriguez et al., 2016).
@@
This trend also explains why the majority (roughly 80%) of the BBH mergers in colossus occur inside the cluster (see Table 1). The GW inspiral time of a BBH scales steeply with semi-major axis ($t_{\rm insp}\propto a^{4}$; Peters, 1964). This means BBHs in more massive clusters that are near the critical orbital separation for dynamical ejection are more susceptible to merging inside their host before ejection can occur (see also, e.g., Rodriguez et al., 2019; Kremer et al., 2020c; Antonini & Gieles, 2020). Indeed, this explains the trend shown in the right-panel of Figure 7 where the ratio of in-cluster to ejected mergers increases with cluster mass.
@@ 5.1 Cluster mapping with distance function
For Virgo GC properties, we use the results from the “ACS Virgo Cluster Survey” (e.g., Côté et al., 2004). In particular, we use the catalog in Jordán et al. (2009) to attain present-day cluster masses and half-light radii.
@@ 5.2 Interpolating the CMC model grid
Figure 11: Estimated total BBH merger populations ($N_{\rm bbh}$, left panel) and generation 2+ merger populations ($N_{\rm 2G+}$, right panel) for each GC in the Virgo cluster, utilizing the scaling relations and weighting scheme described in the text. Constant curves for 1, 10, 100, and 1000 mergers are shown in both panels, obtained by scaling off of a middling CMC model with $\log(M_{\rm cluster}/M_{\odot})=5.7$, $r_{h}=5.0\,\text{pc}$, $N_{\rm bbh}=153$, and $N_{\rm 2G+}=14$ (model 104 in Kremer et al., 2020b). As expected, merger counts generally grow with increasing cluster mass and decreasing $r_{h}$.
@@
Using the weighting scheme of Equation 10, a CMC model that is twice as far from the target Virgo GC as the closest model will be weighted half as much, thrice as far will be weighted one-third as much, etc. An equivalent method can be used to calculate $N_{\rm 2G+}^{\rm Virgo}$, simply by replacing $N_{n}^{\rm bbh}$ with $N_{n}^{\rm 2G+}$ (obtained using the appropriate 2G+ scaling relations). Since Equation 9 may break down when extrapolating over vast mass or $r_{h}$ ranges, only the five closest CMC models are considered, and Virgo GCs which exceed the initial mass of colossus are excluded from the rest of the analysis ($\log M_{\rm cl}>6.8$, only six GCs in the Jordán et al. (2009) catalog).

### v2

@@ 3.2 Black hole mergers
The $243.7~M_{\odot}$ fifth-generation black hole is particularly noted for its similarity in mass to the recent event GW231123, which has total mass $225^{+26}_{-43}~M_{\odot}$ (GW231123). Recent work in PassengerGW23hierarchical placed low probability on a 2G+2G merger being responsible for GW231123, and we find this conclusion consistent with both colossus and lower-mass models. The highest-mass 2G+2G remnant in colossus, being $153.5~M_{\odot}$, lies well below the total mass of GW231123. On the other hand, the 3G+ mergers exclusive to GCs as massive as colossus are more comparable to the mass of GW231123. This suggests that the most massive GCs (comparable to masses of nuclear star clusters) indeed provide a channel for high-mass GW observations, via 3G+ mergers (e.g. AntoniniRasio2016; FragioneRasio2023; GW231123). However, from the perspective of spins, we find the highest-mass mergers formed via 3G+ mergers in colossus generally have spin values lower than the $\chi_{1}=0.90^{+0.10}_{-0.19}$ and $\chi_{2}=0.80^{+0.20}_{-0.51}$ values quoted for GW231123 at $90\%$ confidence (for reasons described in Figure 4). In this case, spins may pose a challenge for formation of GW231123-like events via 3G+ mergers (for further discussion, see PassengerGW23hierarchical). Alternative formation channels that may produce high spins include accretion in active galactic nuclei disks (e.g., Tagawa2020; Vajpeyi2022; Bartos2025) or from stars in dense clusters (e.g., Kiroglu2025c), chemically homogeneous evolution (e.g, Marchant2016; Marchant2024; Stegmann2025), and tidal spin up in stellar binaries (e.g., Bavera2021; Ma2023; Qin2023).
@@
In the left panel of Figure 6 we show the delay time distribution for all mergers in colossus compared to low-mass CMC models (see Figure 2 and Section 4). As shown, more massive clusters generally feature longer delay times, simply because more massive clusters have longer relaxation times ($t_{\rm rlx}\propto M_{\rm cl}^{1/2}$; e.g., BinneyTremaine2008). However at late times ($t\gtrsim 8\,$Gyr), we see the delay time distribution for colossus steepens relative to lower-mass models. This occurs because unlike lower-mass models, nearly all of the mergers in colossus occur inside the cluster, which in general have shorter GW inspiral times relative to ejected mergers (e.g., Rodriguez2018; Kremer2026_review). We address this point further in Section 4.
In Table 1, we list the number of BBH mergers of various types in the colossus model. “Ejected” mergers refer to BBHs that merge following dynamical ejection from the cluster, “2-body insp.” mergers refer to BBHs that merge inside their host cluster following dynamical hardening and GW inspiral, and “2-body cap.”, “3-body cap.”, and “4-body cap.” refer to in-cluster mergers that occur via GW capture during single-single, binary-single, and binary-binary encounters respectively. For further discussion of the distinction between these various channels, see kremer2020modeling.
@@ 3.3 Colossus versus Behemoth
The most comparable CMC model run to-date is the behemoth model of Rodriguez2020_SSC, which was computed as part of the “Great Balls of FIRE” project designed to perform cluster simulations with initial conditions and time-dependent tidal forces extracted from a cosmological simulation (Grudic2023; Rodriguez2023; Bruel2024A). The behemoth model also contained roughly $10^{7}$ stars at birth, but featured two key differences relative to our colossus model: (1) behemoth was run for only 44\,Gyr before it was fully disrupted by external tidal forces in its host galaxy and (2) behemoth is higher metallicity (roughly $0.5\,Z_{\odot}$), which translates to much lower mass black holes due to the influence of metallicity-dependent stellar mass loss due to winds (Vink2001). For example, the median primary mass for all first-generation black hole mergers in behemoth is $16.7\,M_{\odot}$, while it is $30.3\,M_{\odot}$ for colossus. From a numerical perspective, clusters with higher-mass black holes are computationally slower as the black holes undergo deeper collapse episodes driven by enhanced mass segregation (e.g., kremer2020modeling; Rodriguez2022). This in turn leads to shorter time steps compared to higher-metallicity simulations, which results in longer run times. Nonetheless, the behemoth model features similar conclusions to those drawn from our new colossus model, namely that the vast majority of BBH mergers occur inside the cluster and that extended chains of hierarchical mergers are enabled by the relatively higher cluster escape velocity (for further detail and discussion of comparison to specific LVK events, see Rodriguez2020_SSC).
@@ 3.4 Other transient events
Although we focus here primarily on BBH mergers, for completeness, we also include in Table 1 the number of other types of mergers and collisions occuring in the colossus simulation. Although roughly 6,0006{,}000 neutron stars are formed and retained in the cluster via stellar evolution, zero binary neutron star mergers occur throughout the simulation. This is consistent with findings of previous studies (notably Ye2019) that have shown binary neutron star mergers are quite rare in clusters that have not yet reached cluster core collapse. We identify $404$ total white dwarf binary mergers in the simulation, all of which merge inside the cluster. This is also consistent with expectations from previous studies (e.g., Kremer2021_wd) that have shown (1) white dwarf mergers preferentially occur inside their host cluster relative to BBH mergers, owing to their relatively small masses and correspondingly small semi-major axis values required for dynamical ejection and (2) similar to neutron star mergers, white dwarf mergers are rare relative to BH mergers in non-core-collapsed clusters.
We also list at the bottom of Table 1 counts of other types of collisions. Notably, we find $759$ BH+star collisions occur over in the complete simulation. For comparison, $85$ and $37$ BH+star collisions occur in the corresponding CMC Cluster Catalog models with $N=32\times 10^{5},r_{v}=2\,$pc and $N=16\times 10^{5},r_{v}=2\,$pc, respectively. These BH+star collisions may give rise to bright electromagnetic transients powered by accretion of the disrupted star onto the BH (Kremer2019_tde; Kremer2022_sph; Kremer2023_tde; Kiroglu2023), potentially similar to transients events observed recently in the outskirts of nearby old massive elliptical galaxies (Nicholl2023). Similar to BBH mergers, we find there is a strong preference for BH+star collisions to occur in the most massive GCs, which may aid in potential follow-up of observed events as more massive GCs are brighter.
@@ 4.1 Scaling of BBH mergers with cluster mass
For comparison to colossus, we utilize the same five lower-mass CMC Cluster Catalog models plotted in previous figures, with the same initial conditions except for initial cluster $N$. In Figure 7, we plot the number of BBH mergers in all six models versus initial cluster mass, $M_{\rm cl,0}$, and initial $N$, sorted by merger channel, hierarchical generation, and location. In the left panel of Figure 7, we plot the total number of BBH mergers as a solid black curve, alongside three separate curves for the distinct merger channels (for review, see Kremer2026_review): ejected mergers (blue diamonds), in-cluster 2-body inspirals (orange squares), and in-cluster GW capture mergers (red triangles). Note that GW captures here combines single-single, binary-single, and binary-binary capture channels listed in Table 1. A linear regression fit to the total BBH curve in log-log space yields the best-fit relation
@@
For comparison, AntoniniGieles2020 uses semi-analytic methods to find $N_{\rm bbh}\propto M_{\rm cl,0}^{1.6}$ for BBH mergers occurring at late times (t>8t>8\,Gyr). Hong2018 used Monte Carlo models performed with the MOCCA code (Giersz2013) to find $N_{\rm bbh}\propto M_{\rm cl,0}^{1.3}$ for mergers occurring across all times. The $N_{\rm bbh}\propto M_{\rm cl,0}^{1.42}$ scaling found here is roughly consistent with these previous studies. We also performed a similar regression for present-day cluster mass $M_{\rm cl,f}$ (see Figure 2) and found a scaling of $N_{\rm bbh}\propto M_{\rm cl,f}^{1.36}$. Since the scaling with initial and final cluster masses are very similar, henceforth we simply adopt $N_{\rm bbh}\propto M_{\rm cl}^{1.4}$. Regression fits to other CMC models with varying $r_{v}$ show this scaling relation remains consistent, independent of $r_{v}$.
@@ 4.2 Scaling of BBH mergers with half-light radius
Similar to cluster mass, BBH merger populations are also expected to vary with cluster radius (e.g. kremer2020modeling). We utilize 80 CMC models of varying initial $N$ and $r_{v}$ to derive a scaling relation in present-day cluster $r_{h}$, but exclude colossus in this analysis since no other model has comparable mass.
@@
As motivated by previous studies (Hong2018; AntoniniGieles2020), we initially attempted to fit a power-law correlation to $r_{h}$ of the form $N_{\rm bbh}\propto r_{h}^{-\alpha}$, but found that power-law fits were unable to produce accurate results for the $r_{v}=0.5~\text{pc}$ CMC models (previous studies did not consider $r_{v}<1~\text{pc}$ clusters in their analyses). We were able to reproduce AntoniniGieles2020’s $N_{\rm bbh}\propto r_{h}^{-0.9}$ relation by excluding the $r_{v}=0.5~\text{pc}$ models from the regression fit, yielding $N_{\rm bbh}\propto r_{h}^{-0.97}$ (plotted as a dashed curved in Figure 8), but this fit significantly overestimates $N_{\rm bbh}$ at low $r_{h}$. We instead utilize an exponential function of the form $N_{\rm bbh}\propto e^{-\alpha r_{h}}$ to better fit the data. The best-fit curves for the total BBH merger count with varying $r_{h}$ obtained through exponential regression are given in Equation 5, and plotted in Figure 8. The $N=2$ and $N=4$ models did not exhibit sufficiently distinct BBH merger populations, so they were grouped into one curve.
@@ 4.3 Ejected BBH properties
In the dense environment of GCs, BBHs experience repeated close encounters with other single stars and binary systems. Each of these encounters transfer energy from the BBH’s internal energy into the kinetic energy of the interacting objects, which leads to BBH hardening and dynamical recoil (e.g., Rodriguez2016mergers; Kremer2026_review). Energetics arguments show center-of-mass recoil velocity of a binary is comparable to its orbital velocity $v_{\rm orb}\approx\sqrt{Gm_{\rm bh}/a}$. Therefore, as a BBH hardens, it recoils more. The characteristic minimum separation a binary can achieve via dynamical hardening is set by its host cluster’s escape velocity (Equation 1). Once a BBH becomes sufficiently compact, its recoil velocity will exceed its host’s escape velocity, and it will be ejected from the cluster. A subset of these ejected binaries will have sufficiently short GW inspiral times to ultimately merge within a Hubble time. Since higher-mass GCs have higher escape energy thresholds, ejected BBH systems of a higher-mass GC will have higher binding energies, shorter semi-major axes, and correspondingly shorter GW inspiral times (Rodriguez2016mergers).
@@
This trend also explains why the majority (roughly 80%) of the BBH mergers in colossus occur inside the cluster (see Table 1). The GW inspiral time of a BBH scales steeply with semi-major axis ($t_{\rm insp}\propto a^{4}$; Peters1964). This means BBHs in more massive clusters that are near the critical orbital separation for dynamical ejection are more susceptible to merging inside their host before ejection can occur (see also, e.g., Rodriguez2019; Kremer2020; AntoniniGieles2020). Indeed, this explains the trend shown in the right-panel of Figure 7 where the ratio of in-cluster to ejected mergers increases with cluster mass.
@@ 5.1 Cluster mapping with distance function
For Virgo GC properties, we use the results from the “ACS Virgo Cluster Survey” (e.g., Cote2004). In particular, we use the catalog in Jordan2009 to attain present-day cluster masses and half-light radii.
Figure 11: Estimated total BBH merger populations ($N_{\rm bbh}$, left panel) and generation 2+ merger populations ($N_{\rm 2G+}$, right panel) for each GC in the Virgo cluster, utilizing the scaling relations and weighting scheme described in the text. Constant curves for 1, 10, 100, and 1000 mergers are shown in both panels, obtained by scaling off of a middling CMC model with $\log(M_{\rm cluster}/M_{\odot})=5.7$, $r_{h}=5.0\,\text{pc}$, $N_{\rm bbh}=153$, and $N_{\rm 2G+}=14$ (model 104 in kremer2020modeling). As expected, merger counts generally grow with increasing cluster mass and decreasing $r_{h}$.
@@ 5.2 Interpolating the CMC model grid
Using the weighting scheme of Equation 10, a CMC model that is twice as far from the target Virgo GC as the closest model will be weighted half as much, thrice as far will be weighted one-third as much, etc. An equivalent method can be used to calculate $N_{\rm 2G+}^{\rm Virgo}$, simply by replacing $N_{n}^{\rm bbh}$ with $N_{n}^{\rm 2G+}$ (obtained using the appropriate 2G+ scaling relations). Since Equation 9 may break down when extrapolating over vast mass or $r_{h}$ ranges, only the five closest CMC models are considered, and Virgo GCs which exceed the initial mass of colossus are excluded from the rest of the analysis ($\log M_{\rm cl}>6.8$, only six GCs in the Jordan2009 catalog).